# Learning Visual Hierarchies in Hyperbolic Space for Image Retrieval

### v3

@@ 1 Introduction
Humans organize knowledge of the world into hierarchies [27] for efficient knowledge management. Developing models that encode such hierarchies is crucial for creating systems with holistic world understanding aligned with human perception. While this topic spans various modalities, we focus on encoding hierarchies in the visual domain. For many large image datasets, objects can be organized according to latent hierarchies [27], as evidenced by power law distributions [35]. However, most prevalent image understanding models [37, 17, 15, 5, 16] focus on preserving visual similarity, and consequently, learning hierarchies in the visual domain remains relatively unexplored. These limitations hinder models’ ability to generalize to tasks requiring hierarchical reasoning such as understanding scenes, object parts, and their interactions at multiple levels of abstraction. We illustrate in Fig. 1 an example that shows the complexity of a visual hierarchy where elements share similarities both within and across categories.
@@
Recent works have demonstrated the utility of hyperbolic representation spaces for capturing hierarchical relationships in an unsupervised setting [7, 33, 31]. This emergent latent structure is appealing; nevertheless, meaningful hierarchies are often task and data dependent, and aligning model behavior with such human-defined hierarchies is essential for many applications. To this end, we introduce a learning paradigm that can encode multi-level hierarchies as entailment pairs in hyperbolic space. As a concrete example, we first define a general part-based image hierarchy using object and part level classification annotations within and across images. Then, we introduce a model capable of structuring the latent space to preserve the defined hierarchy using only image/object level information. To our knowledge, we are the first to encode multi-level complex visual hierarchies without relying on explicit hierarchical labels or additional modalities. Finally, we introduce an evaluation metric to effectively measure hierarchical image retrieval.
@@
Note that, hierarchy is an asymmetric relationship and has a high branching factor (see Fig. 1). To this end, we adopt the hyperbolic geometry as it provides a continuous approximation of such tree-like structures [27, 36]. To enforce the hierarchy, we break it into pairwise entailment relationships between images, objects, and parts, at multiple levels within an image as well as across images at category level. For pairwise entailment (asymmetric), we adapt the recently proposed angle-based asymmetric distance metric [33] within the contrastive learning paradigm, and extend it to handle cases with multiple positive relationships. In contrast to symmetric distances such as the inner-product used in [28, 21, 12], this angle-based distance offers an additional degree of freedom along the radial axis to form emergent structures in the latent space.
@@
For experimentation, we build a dataset of visual hierarchies using the bounding box annotations of OpenImages [24]. Our dataset includes entailment relationships between scenes, objects, and parts, within a single image as well as across images at the category level. Similarly, for hierarchical retrieval evaluation, we use the full training set to create ground truth hierarchy trees per scene/object. Additionally, we design a metric for evaluating hierarchical retrieval based on the optimal transport distance between the label distribution of the retrieval set and ground truth label distribution within the hierarchy tree. Combined with Recall@k metrics, this demonstrates that our method captures semantic and structural information, transcending mere visual similarity. Furthermore, to the best of our knowledge, our model is the first to generalize to out-of-domain image hierarchies, achieving strong performance on unseen and out-of-domain datasets.
@@ 2 Preliminaries
We briefly review essential concepts in hyperbolic geometry here. We refer the reader to  [34] for a comprehensive treatment. Hyperbolic spaces are Riemannian manifolds with constant negative curvature and are fundamentally different from Euclidean or spherical space which has zero or constant positive curvature, respectively. The negative curvature enables properties such as divergence of parallel lines and exponential volume growth with radius [3]. This volume growth property makes hyperbolic space an ideal candidate for embedding hierarchical and graph structured data [27], and has found many machine learning applications.
@@ 3.1 Part-Based Image Hierarchy
Visual hierarchies can be established in different ways, depending on the application. In this work, we are interested in a hierarchy that encapsulates the semantic relationship among objects in a scene. For this, scene-object-part hierarchy is appealing as it is useful for applications such as fine-grained object retrieval, object localization, and general scene understanding. This hierarchy has also been shown to emerge in hyperbolic image embeddings [21, 33].
@@ 3.2 Angle-Based Entailment Loss
We require an asymmetric distance function to enforce pairwise entailment relationships in hyperbolic space. To this end, we adapt the recently proposed hyperbolic-angle-based entailment loss [33], a smooth contrastive variant of the entailment cone loss [9]. The angle-only loss, without distance constraints on image pairs, provides flexibility to be distributed along the radial axis, allowing embeddings to align with the tree structure while preserving pairwise entailment.
Our loss is a bidirectional version of [33], as illustrated graphically in Fig. 3. In particular, given embeddings of an entailment pair, $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$ in the tangent space, where $\mathbf{x}$ entails $\mathbf{y}$, we maximize the angles $\beta_{1}$ and $\alpha_{2}$. This enforces entailment in a bidirectional manner. The angles $\beta_{1}$ and $\alpha_{2}$ can be computed using the exterior angle as follows:
@@
where $\operatorname{ext}(\mathbf{x},\mathbf{y})$ is the exterior angle between $\mathbf{x}$ and $\mathbf{y}$ and takes the following form [7]:
@@
In contrast to [33], in our case an embedding can belong to multiple entailment pairs in a batch. This corresponds to a case of multiple positives in the contrastive loss. Thus, we employ the InfoNCE loss [29] to align all entailment pairs while pushing apart the rest of the negative pairs. Precisely, let $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}$ be a batch of entailment pairs, then the InfoNCE loss for parent-to-child can be written as:
$L^{p\to c}(\mathcal{D},\kappa)=$ (8)
$-\mathbb{E}_{\mathcal{D}}\!\left[\log\frac{\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i})}{\tau}\right)}{\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i})}{\tau}\right)+\sum_{\mathbf{y}^{-}\in\mathcal{N}_{i}}\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i}^{-})}{\tau}\right)}\right]\ ,$
where $\mathcal{N}_{i}$ denotes the set of samples that do not have an entailment relationship with the parent embedding $\mathbf{x}_{i}$. Here, $\kappa:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the similarity function, and $\tau$ is a learnable temperature parameter initialized to 0.07 following [7]. Now, our bidirectional entailment loss can be written as:
@@
In our implementation, we use a shared image encoder for both parent and child embeddings. Following the reparametrization of [7], we encode the space component of the Lorentz model in the tangent space at origin and project it onto the hyperboloid using the exponential map, enabling contrastive entailment angle loss computation in hyperbolic space.
@@ 3.3 Hierarchical Retrieval Evaluation
To this end, consider the parent-to-child relationship, and let $\mathcal{H}_{I}$ denote the hierarchy tree originating from the query image $I$, containing $m$ labels. Then, the labels in $\mathcal{H}_{I}$ are the ground truth labels for the hierarchical retrieval for image $I$. Now, let $\mathbf{h}_{I}\in\mathbf{R}^{m}$ be the precomputed label distribution of $\mathcal{H}_{I}$. Then, our optimal transport (1-D Wasserstein) distance between the retrieved label distribution $\mathbf{r}_{I}$ and $\mathbf{h}_{I}$ is:
@@ 4 Related Work
Hyperbolic geometry allows for exponential volume growth with respect to the radius [3], making it effective for embedding hierarchical structures. This advantage has led to significant research into leveraging hyperbolic representations for various data types, including molecular structures [42], 3D [4, 18, 40], images [21, 13, 12, 22, 31], text data [38, 38, 10, 43, 19], and vision-language data [7, 33, 23, 2].
Hyperbolic embeddings can be learned through standard deep learning layers [20] with hyperbolic projection [27] or using hyperbolic neural networks [10]. Many prior NLP, computer vision and knowledge graph studies learn hierarchies from partially order data [10, 39, 25], or minimizing geodesic distance or maximizing similarities [27, 28, 41, 12]. Ganea et al. [10] introduced an angle-based entailment cone loss which pushes child nodes into the cone emanating from the parent node embedding. This approach has been applied to both text [10] and image data [8] with label hierarchies. Recently, this hyperbolic entailment loss was adapted for contrastive learning to develop representations in vision-language models [7, 33, 2, 30]. However, such methods remain relatively unexplored in the image domain. We adapted the angle-based entailment loss from [33] to encode part-based image hierarchies. Many previous works are limited to learning hierarchies for single-class images using predefined labeled hierarchies, such as ImageNet [26, 21] or hand-labeled data [8]. In this work, we propose a learning method that fine-tunes pre-trained models on large-scale datasets for general images (with multiple classes per image) without hierarchical labels. The most relevant approach, HCL [12], models simple scene-to-object hierarchies, whereas we capture more complex, multi-level part-based hierarchies directly from image data, extending beyond visual similarity.
@@ 5 Experiments
For training and evaluation, we construct HierOpenImages, a novel dataset containing pairwise part-based image hierarchies built from the OpenImages dataset [24]. We further evaluate generalization on out-of-domain datasets and hierarchies.
@@
We evaluate the performance of learning multi-level image hierarchies using two popular visual encoders 1) CLIP ViT (B/16) [32], pretrained on large-scale image-text pairs from the Internet, and 2) MoCo-v2 (ResNet-50) [6], pretrained on ImageNet. Note that both CLIP and MoCo-v2 models are fine-tuned and evaluated in an image encoder only setting. We use these pretrained image models as our baseline and compare our proposed angle-based hyperbolic method against its Euclidean counterpart. Each model is fine-tuned for a single epoch on HierOpenImages using the proposed contrastive angle-based entailment loss, and $\dagger$ denotes fine-tuning. The retrieval distance function (Dist. Func.) aligns with the scoring function used during training. For example, minimizing hyperbolic angles (Hyp Ang.) for CLIP-hyp and Euclidean angles (Euc Ang.) for CLIP-euc model.
We also compare the state-of-the-art hyperbolic hierarchical image-only model HCL [12], which is trained on scene-to-object relationships from the OpenImages dataset [24] in both hyperbolic and Euclidean space. Accordingly, we evaluate the retrieval performance of HCL [12] using both hyperbolic distance and cosine similarity. For completeness, we also fine-tuned HCL [12] on the same HierOpenImages dataset and included it in our comparisons. Note that fine-tuning or evaluating with a text encoder is not possible on the current HierOpenImages dataset. Furthermore, previous image-only supervised methods, trained on predefined single-class image hierarchies [8, 21] are not directly comparable. These methods require all image classes to be predefined during training, which is not feasible for complex multi-object scenes in the OpenImages [24] and LVIS [14] datasets, where images contain multiple objects with diverse class labels.
@@ 5.1 Main Results
Denoting the full image as parent and the bounding box as child, we evaluate retrieval tasks in both child-to-parent and parent-to-child directions. Table 1 shows the retrieval accuracy of $\text{top-k}=\{5,10,50,100\}$. We notice a significant and consistent improvement with our proposed hyperbolic model, across all metrics and model variants. This highlights the relevance of our angle-based entailment loss and the advantages of learning hierarchical image embeddings in hyperbolic space. While for HCL [12], only a slight performance increase was observed after fine-tuning.
@@
CLIP ViT CLIP Cos Sim. 26.73 25.95 23.69 22.70
CLIP-euc Euc Ang. 67.83 68.37 67.12 66.04
CLIP-hyp Hyp Ang. 73.37 72.59 69.84 68.62
HCL Cos Sim. 15.11 14.49 14.69 14.60
HCL Hyp Dist. 14.46 14.34 14.04 13.92
HCL Hyp Dist. 14.54 14.43 14.16 14.01
MoCo-v2 MoCo Cos Sim. 17.23 16.86 16.33 16.07
MoCo-euc Euc Ang. 54.51 54.16 51.56 49.53
MoCo-hyp Hyp Ang. 55.53 55.31 51.53 49.51
@@
CLIP ViT CLIP Cos Sim. 47.52 46.60 43.50 42.08
CLIP-euc Euc Ang. 65.38 65.70 66.01 65.79
CLIP-hyp Hyp Ang. 66.02 66.63 66.50 65.91
HCL Cos Sim. 16.08 15.93 15.84 15.74
HCL Hyp Dist. 16.16 16.04 15.60 15.45
HCL Hyp Dist. 17.07 16.57 15.97 15.69
MoCo-v2 MoCo Cos Sim. 18.49 18.37 17.73 17.45
MoCo-euc Euc Ang. 47.11 46.86 46.95 47.22
MoCo-hyp Hyp Ang. 52.01 51.64 50.83 50.51
@@
Recall $\uparrow$ CLIP Cos Sim. 51.33 64.28 77.02
CLIP-euc Euc Ang. 72.96 80.98 87.97
CLIP-hyp Hyp Ang. 74.06 82.34 88.79
OT Distance $\downarrow$ CLIP Cos Sim. 23.81 24.60 25.03
CLIP-euc Euc Ang. 17.00 20.00 22.45
CLIP-hyp Hyp Ang. 16.36 19.27 22.21
@@
Recall $\uparrow$ HCL Cos Sim. 46.05 61.24 76.44
HCL Hyp Dist. 46.06 61.14 76.16
HCL Hyp Dist. 45.81 60.77 75.77
MoCo Cos Sim. 48.38 63.63 77.89
MoCo-euc Euc Ang. 55.81 71.86 83.42
MoCo-hyp Hyp Ang. 56.09 71.61 83.13
OT Distance $\downarrow$ HCL Cos Sim. 24.23 24.67 25.09
HCL Hyp Dist. 25.72 25.98 26.04
HCL Hyp Dist. 26.12 26.32 26.28
MoCo Cos Sim. 24.05 24.44 24.93
MoCo-euc Euc Ang. 13.72 16.83 20.17
MoCo-hyp Hyp Ang. 12.41 15.29 19.09
@@
Figure 4: Precision-Recall curves of CLIP ViT models on hierarchical retrieval. Dotted lines show models trained only on hierarchical entailment data within the same images; solid lines represent models trained with additional cross-image scene-to-object samples. Angle or cosine similarity threshold values are marked by text.
@@
As shown in Table 2, our hyperbolic model better captures the hierarchical distribution of the test set compared to the Euclidean model, achieving better OT distance in all cases, and in 4 out of 6 instances for the recall.
All fine-tuned models using our angle-based entailment loss show significant improvement over the baseline models. Notably, even after fine-tuning, HCL [12] shows a decline in hierarchical retrieval performance, indicating its difficulty in learning the complex visual hierarchies in the training data.
@@
Fig. 4 shows the Precision-Recall (PR) curve with increasing $\beta_{1}$ angle thresholds (0 to $\pi$) for CLIP-hyp and CLIP-euc, and with cosine similarity thresholds (0 to 1) for the baseline CLIP. The proposed hyperbolic model trained without cross-image samples substantially outperforms the Euclidean model and even exceeds the Euclidean model with cross-image samples, indicating it implicitly learns the underlying structure. When cross-image samples are introduced, the hyperbolic model still leads, though the performance gap narrows.
@@
Following [7, 11, 33], we use parent-to-child image traversals with results ordered by increasing embedding norm, to illustrate the latent space, which: (1) aligns with the predefined scene-object-part hierarchy, placing high-level objects near the origin. (2). groups diverse objects under the same lower hierarchical branch, even if they are visually distinct (e.g., keyboard and monitor under the studio image). Moreover, Figure 10 shows an example of the model’s emergent structure where high-frequency, ambiguous image crops are naturally pushed toward the boundary of hyperbolic space. Find details in supplementary materials.

### v4

@@ 1 Introduction
Humans organize knowledge of the world into hierarchies [nickel2017poincare] for efficient knowledge management. Developing models that encode such hierarchies is crucial for creating systems with holistic world understanding aligned with human perception. While this topic spans various modalities, we focus on encoding hierarchies in the visual domain. For many large image datasets, objects can be organized according to latent hierarchies [nickel2017poincare], as evidenced by power law distributions [ravasz2003hierarchical]. However, most prevalent image understanding models [schroff2015facenet, hoffer2015deep, he2016deep, chen2020simple, he2020momentum] focus on preserving visual similarity, and consequently, learning hierarchies in the visual domain remains relatively unexplored. These limitations hinder models’ ability to generalize to tasks requiring hierarchical reasoning such as understanding scenes, object parts, and their interactions at multiple levels of abstraction. We illustrate in Fig. 1 an example that shows the complexity of a visual hierarchy where elements share similarities both within and across categories.
@@
Recent works have demonstrated the utility of hyperbolic representation spaces for capturing hierarchical relationships in an unsupervised setting [desai2023hyperbolic, ramasinghe2024accept, qiu2024hihpq]. This emergent latent structure is appealing; nevertheless, meaningful hierarchies are often task and data dependent, and aligning model behavior with such human-defined hierarchies is essential for many applications. To this end, we introduce a learning paradigm that can encode multi-level hierarchies as entailment pairs in hyperbolic space. As a concrete example, we first define a general part-based image hierarchy using object and part level classification annotations within and across images. Then, we introduce a model capable of structuring the latent space to preserve the defined hierarchy using only image/object level information. To our knowledge, we are the first to encode multi-level complex visual hierarchies without relying on explicit hierarchical labels or additional modalities. Finally, we introduce an evaluation metric to effectively measure hierarchical image retrieval.
@@
Note that, hierarchy is an asymmetric relationship and has a high branching factor (see Fig. 1). To this end, we adopt the hyperbolic geometry as it provides a continuous approximation of such tree-like structures [nickel2017poincare, sala2018representation]. To enforce the hierarchy, we break it into pairwise entailment relationships between images, objects, and parts, at multiple levels within an image as well as across images at category level. For pairwise entailment (asymmetric), we adapt the recently proposed angle-based asymmetric distance metric [ramasinghe2024accept] within the contrastive learning paradigm, and extend it to handle cases with multiple positive relationships. In contrast to symmetric distances such as the inner-product used in [nickel2018learning, khrulkov2020hyperbolic, ge2023hyperbolic], this angle-based distance offers an additional degree of freedom along the radial axis to form emergent structures in the latent space.
@@
For experimentation, we build a dataset of visual hierarchies using the bounding box annotations of OpenImages [kuznetsova2020open]. Our dataset includes entailment relationships between scenes, objects, and parts, within a single image as well as across images at the category level. Similarly, for hierarchical retrieval evaluation, we use the full training set to create ground truth hierarchy trees per scene/object. Additionally, we design a metric for evaluating hierarchical retrieval based on the optimal transport distance between the label distribution of the retrieval set and ground truth label distribution within the hierarchy tree. Combined with Recall@k metrics, this demonstrates that our method captures semantic and structural information, transcending mere visual similarity. Furthermore, to the best of our knowledge, our model is the first to generalize to out-of-domain image hierarchies, achieving strong performance on unseen and out-of-domain datasets.
@@ 2 Preliminaries
We briefly review essential concepts in hyperbolic geometry here. We refer the reader to  [ratcliffe1994foundations] for a comprehensive treatment. Hyperbolic spaces are Riemannian manifolds with constant negative curvature and are fundamentally different from Euclidean or spherical space which has zero or constant positive curvature, respectively. The negative curvature enables properties such as divergence of parallel lines and exponential volume growth with radius [bridson2013metric]. This volume growth property makes hyperbolic space an ideal candidate for embedding hierarchical and graph structured data [nickel2017poincare], and has found many machine learning applications.
@@ 3.1 Part-Based Image Hierarchy
Visual hierarchies can be established in different ways, depending on the application. In this work, we are interested in a hierarchy that encapsulates the semantic relationship among objects in a scene. For this, scene-object-part hierarchy is appealing as it is useful for applications such as fine-grained object retrieval, object localization, and general scene understanding. This hierarchy has also been shown to emerge in hyperbolic image embeddings [khrulkov2020hyperbolic, ramasinghe2024accept].
@@ 3.2 Angle-Based Entailment Loss
We require an asymmetric distance function to enforce pairwise entailment relationships in hyperbolic space. To this end, we adapt the recently proposed hyperbolic-angle-based entailment loss [ramasinghe2024accept], a smooth contrastive variant of the entailment cone loss [ganea2018entail]. The angle-only loss, without distance constraints on image pairs, provides flexibility to be distributed along the radial axis, allowing embeddings to align with the tree structure while preserving pairwise entailment.
Our loss is a bidirectional version of [ramasinghe2024accept], as illustrated graphically in Fig. 3. In particular, given embeddings of an entailment pair, $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$ in the tangent space, where $\mathbf{x}$ entails $\mathbf{y}$, we maximize the angles $\beta_{1}$ and $\alpha_{2}$. This enforces entailment in a bidirectional manner. The angles $\beta_{1}$ and $\alpha_{2}$ can be computed using the exterior angle as follows:
@@
where $\operatorname{ext}(\mathbf{x},\mathbf{y})$ is the exterior angle between $\mathbf{x}$ and $\mathbf{y}$ and takes the following form [desai2023hyperbolic]:
@@
In contrast to [ramasinghe2024accept], in our case an embedding can belong to multiple entailment pairs in a batch. This corresponds to a case of multiple positives in the contrastive loss. Thus, we employ a multi-positive variant of the InfoNCE loss [oord2018representation] to align all entailment pairs while pushing apart the rest of the negative pairs. Precisely, let $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}$ be a batch of entailment pairs. For each parent embedding $\mathbf{x}_{i}$, let $\mathcal{P}_{i}$ denote the set of child embeddings in the batch that have an entailment relationship with $\mathbf{x}_{i}$, and let $\mathcal{A}$ denote the set of all candidate embeddings in the batch. Then the InfoNCE loss for parent-to-child can be written as:
$L^{p\to c}(\mathcal{D},\kappa)=-\mathbb{E}_{\mathcal{D}}\!\left[\log\frac{\sum\limits_{\mathbf{y}\in\mathcal{P}_{i}}\exp\!\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y})}{\tau}\right)}{\sum\limits_{\mathbf{y}\in\mathcal{A}}\exp\!\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y})}{\tau}\right)}\right]\ ,$ (8)
where $\kappa:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the similarity function, and $\tau$ is a learnable temperature parameter initialized to 0.07 following [desai2023hyperbolic]. Now, our bidirectional entailment loss can be written as:
@@
In our implementation, we use a shared image encoder for both parent and child embeddings. Following the reparametrization of [desai2023hyperbolic], we encode the space component of the Lorentz model in the tangent space at origin and project it onto the hyperboloid using the exponential map, enabling contrastive entailment angle loss computation in hyperbolic space.
@@ 3.3 Hierarchical Retrieval Evaluation
Consider the parent-to-child relationship, and let $\mathcal{H}_{I}$ denote the hierarchy tree originating from the query image $I$, containing $m$ labels. Then, the labels in $\mathcal{H}_{I}$ are the ground truth labels for the hierarchical retrieval for image $I$. Now, let $\mathbf{h}_{I}\in\mathbf{R}^{m}$ be the precomputed label distribution of $\mathcal{H}_{I}$. Then, our optimal transport (1-D Wasserstein) distance between the retrieved label distribution $\mathbf{r}_{I}$ and $\mathbf{h}_{I}$ is:
@@ 4 Related Work
Hyperbolic geometry allows for exponential volume growth with respect to the radius [bridson2013metric], making it effective for embedding hierarchical structures. This advantage has led to significant research into leveraging hyperbolic representations for various data types, including molecular structures [yu2022skin], 3D [chen2020learning, hsu2021capturing, wang2023hypliloc], images [khrulkov2020hyperbolic, guo2022clipped, ge2023hyperbolic, kim2023hier, qiu2024hihpq], text data [tifrea2018poincar, tifrea2018poincar, ganea2018hyperbolic, zhu2020hypertext, jeong2024simple], and vision-language data [desai2023hyperbolic, ramasinghe2024accept, kong2024hyperbolic, alper2024hierarcaps].
Hyperbolic embeddings can be learned through standard deep learning layers [khan2022transformers] with hyperbolic projection [nickel2017poincare] or using hyperbolic neural networks [ganea2018hyperbolic]. Many prior NLP, computer vision and knowledge graph studies learn hierarchies from partially order data [ganea2018hyperbolic, vilnis2018probabilistic, liu2019hyperbolic], or minimizing geodesic distance or maximizing similarities [nickel2017poincare, nickel2018learning, yan2021unsupervised, ge2023hyperbolic]. Ganea et al. [ganea2018hyperbolic] introduced an angle-based entailment cone loss which pushes child nodes into the cone emanating from the parent node embedding. This approach has been applied to both text [ganea2018hyperbolic] and image data [dhall2020hierarchical] with label hierarchies. Recently, this hyperbolic entailment loss was adapted for contrastive learning to develop representations in vision-language models [desai2023hyperbolic, ramasinghe2024accept, alper2024hierarcaps, pal2024compositional]. However, such methods remain relatively unexplored in the image domain. We adapted the angle-based entailment loss from [ramasinghe2024accept] to encode part-based image hierarchies. Many previous works are limited to learning hierarchies for single-class images using predefined labeled hierarchies, such as ImageNet [liu2020hyperbolic, khrulkov2020hyperbolic] or hand-labeled data [dhall2020hierarchical]. In this work, we propose a learning method that fine-tunes pre-trained models on large-scale datasets for general images (with multiple classes per image) without hierarchical labels. The most relevant approach, HCL [ge2023hyperbolic], models simple scene-to-object hierarchies, whereas we capture more complex, multi-level part-based hierarchies directly from image data, extending beyond visual similarity.
@@ 5 Experiments
For training and evaluation, we construct HierOpenImages, a novel dataset containing pairwise part-based image hierarchies built from the OpenImages dataset [kuznetsova2020open]. We further evaluate generalization on out-of-domain datasets and hierarchies.
@@
We evaluate the performance of learning multi-level image hierarchies using two popular visual encoders 1) CLIP ViT (B/16) [radford2021learning], pretrained on large-scale image-text pairs from the Internet, and 2) MoCo-v2 (ResNet-50) [chen2020improved], pretrained on ImageNet. Note that both CLIP and MoCo-v2 models are fine-tuned and evaluated in an image encoder only setting. We use these pretrained image models as our baseline and compare our proposed angle-based hyperbolic method against its Euclidean counterpart. Each model is fine-tuned on HierOpenImages using the proposed contrastive angle-based entailment loss, and $\dagger$ denotes fine-tuning. Note that * denotes the gated entailment angle metric, which first uses the entailment angle as a cone-shaped filter and then applies the pretrained metric within the gated set for fine-grained ranking (see Supp. §B.2).
We compare the state-of-the-art hyperbolic hierarchical image-only model HCL [ge2023hyperbolic], which is trained on scene-to-object relationships from the OpenImages dataset [kuznetsova2020open] in both hyperbolic and Euclidean space. Accordingly, we evaluate the retrieval performance of HCL [ge2023hyperbolic] using both hyperbolic distance and cosine similarity. For completeness, we also fine-tuned HCL [ge2023hyperbolic] on the HierOpenImages dataset. Note that fine-tuning or evaluating with a text encoder is not possible on the HierOpenImages dataset. Furthermore, previous image-only supervised methods, trained on predefined single-class image hierarchies [dhall2020hierarchical, khrulkov2020hyperbolic] are not directly comparable. These methods require all image classes to be predefined during training, which is not feasible for complex multi-object scenes in the OpenImages [kuznetsova2020open] and LVIS [gupta2019lvis] datasets, where images contain multiple objects with diverse class labels.  [^4] Updated results after fixing an extra normalization issue in the code which impacted the results of all methods; overall trends and conclusions remain unaffected.
@@ 5.1 Main Results
Denoting the full image as parent and the bounding box as child, we evaluate retrieval tasks in both child-to-parent and parent-to-child directions. Table 1 shows the retrieval accuracy of $\text{top-k}=\{5,10,50,100\}$. We notice a consistent improvement with our proposed hyperbolic model, across all metrics and model variants. This highlights the relevance of our angle-based entailment loss and the advantages of learning hierarchical image embeddings in hyperbolic space.
@@
CLIP ViT CLIP Cos Sim. 53.04 51.69 47.87 45.79
CLIP-euc Euc Ang..* 75.63 74.65 72.25 70.52
CLIP-hyp Hyp Ang..* 77.28 75.91 72.85 70.94
HCL Hyp Dist. 46.57 44.86 39.64 36.61
HCL Cos Sim. 55.48 54.81 52.23 50.67
HCL Hyp Dist. 39.71 37.55 32.13 29.44
HCL Cos Sim. 56.76 55.85 53.61 52.16
MoCo-v2 MoCo Cos Sim. 53.72 52.71 50.41 49.19
MoCo-euc Euc Ang..* 66.78 66.58 65.42 64.40
MoCo-hyp Hyp Ang..* 68.62 67.42 66.21 64.94
@@
CLIP ViT CLIP Cos Sim. 69.94 67.74 62.50 60.21
CLIP-euc Euc Ang..* 73.74 72.57 70.04 68.78
CLIP-hyp Hyp Ang..* 74.48 73.27 70.78 69.46
HCL Hyp Dist. 50.86 49.81 47.34 46.06
HCL Cos Sim. 53.64 53.04 51.34 50.59
HCL Hyp Dist. 53.29 52.39 50.04 48.74
HCL Cos Sim. 54.99 54.36 52.86 52.03
MoCo-v2 MoCo Cos Sim. 63.65 62.21 58.85 57.28
MoCo-euc Euc Ang..* 68.87 67.94 65.64 64.42
MoCo-hyp Hyp Ang..* 69.50 68.14 65.67 64.46
@@
Recall $\uparrow$ CLIP Cos Sim. 66.63 77.15 86.57
CLIP-euc Euc Ang..* 76.46 85.08 91.38
CLIP-hyp Hyp Ang..* 77.00 85.75 91.89
OT Distance $\downarrow$ CLIP Cos Sim. 21.31 22.74 23.79
CLIP-euc Euc Ang..* 15.65 18.15 21.09
CLIP-hyp Hyp Ang..* 14.96 17.57 20.76
@@
Recall $\uparrow$ HCL Hyp Dist. 51.83 64.84 77.66
HCL Cos Sim. 57.05 67.65 78.71
HCL Hyp Dist. 54.51 66.70 78.83
HCL Cos Sim. 60.13 70.26 80.51
MoCo Cos Sim. 65.01 75.85 85.61
MoCo-euc Euc Ang..* 74.84 83.71 90.46
MoCo-hyp Hyp Ang..* 74.85 83.79 90.61
OT Distance $\downarrow$ HCL Hyp Dist. 24.58 24.84 25.01
HCL Cos Sim. 23.58 24.65 25.11
HCL Hyp Dist. 25.30 25.63 25.53
HCL Cos Sim. 21.94 23.58 24.58
MoCo Cos Sim. 18.60 20.79 22.56
MoCo-euc Euc Ang..* 16.65 18.75 21.38
MoCo-hyp Hyp Ang..* 16.59 18.65 21.29
@@
Figure 4: Precision-Recall curves of CLIP ViT models on hierarchical retrieval. Dotted lines show models trained only on hierarchical entailment data within the same images; solid lines represent models trained with additional cross-image scene-to-object samples.
@@
As shown in Table 2, our hyperbolic model better captures the hierarchical distribution of the test set compared to the Euclidean model, achieving better recall and OT distance in all cases. All fine-tuned models using our angle-based entailment loss show clear improvement over the baseline models. In contrast, HCL [ge2023hyperbolic] shows only a small gain after fine-tuning, suggesting that its training method is not well suited for capturing the complex visual hierarchies in our data.
@@
Fig. 4 shows the Precision-Recall (PR) curve with increasing $\beta_{1}$ angle thresholds (0 to $\pi$) for CLIP-hyp and CLIP-euc, and with cosine similarity thresholds (0 to 1) for the baseline CLIP. Fine-tuning in both Euclidean and Hyperbolic spaces substantially improves performance from the pretrained model. Introducing cross-image pairs further improves recall in the mid-to-high recall regions, indicating better alignment across semantically related but visually diverse examples, with only a small precision drop at the top ranks due to the added variability from cross-image relationships.
@@
Following [desai2023hyperbolic, Ge2021RobustCL, ramasinghe2024accept], we use parent-to-child image traversals with results ordered by increasing embedding norm, to illustrate the latent space, which: (1) aligns with the predefined scene-object-part hierarchy, placing high-level objects near the origin. (2). groups diverse objects under the same lower hierarchical branch, even if they are visually distinct (e.g., keyboard and monitor under the studio image). Moreover, Figure 10 shows an example of the model’s emergent structure where high-frequency, ambiguous image crops are naturally pushed toward the boundary of hyperbolic space. Find details in supplementary materials.